the vagabonds were 200 miles apart at 2 p.m. and were headed toward each other. if they met at 6 p.m. and one was traveling 10 miles per hour faster than the other,what was the speed of each vagabond?
Answer by richwmiller(17219)

r4=200
r=50 mph combined speed
2f+10=50
2f=40
f=20
f+10=30
or
f*4+(f+10)*4=200
f*4+f*4+40=200
8f+40=200
8f=160
f=20
f+10=30
check
4*20+4*30=200
80+120=200
200=200